A train travel from city A to city B with a speed of 100km/h another train travel from city B to city A with a speed of 9km/h. distance between the two cities is 475km.what is the distance traveled by the train started from city A when meeting the train arriving from city B?
Answer by stanbon(75887)

A train travel from city A to city B with a speed of 100km/h another train travel from city B to city A with a speed of 90 km/h. distance between the two cities is 475km.what is the distance traveled by the train started from city A when meeting the train arriving from city B?
----
A to B DATA:
rate = 100 km/hr ; time = t hrs ; Dist = r*t = 100t km
----
B to A DATA:
rate = 90 km/hr ; time = t hrs ; Dist = 90t km
----
Equation:
dist + dist = 475 km
190t = 475
time = 2.5 hrs
----
Ans: dist to meet for A to B train :: 2.5*100 = 250 miles
--------------
Cheers,
Stan H.